# On finding optimal collective variables for complex systems by minimizing the deviation between effective and full dynamics

### v1

@@ 1 Introduction
A commonly adopted strategy is to utilize the fact that the dynamics of high-dimensional systems, e.g. molecular systems in MD and materials science, can often be well characterized using only a few observables, often called collective variables (CVs) or reaction coordinates, of the systems. In particular, many enhanced sampling methods (see [16] for a review) utilize the knowledge about the system’s CVs to perform accelerated sampling and efficient free energy calculations [28], whereas various methods for model reduction rely on this knowledge to construct simpler (lower-dimensional) surrogate models [34, 24, 1]. While these methods have proven to be greatly helpful in understanding dynamical behaviors of complex systems, the accuracy of estimated quantities and surrogate models offered by these methods crucially depend on the choices of CVs. Given a set of CVs, analyzing the accuracy of surrogate models in approximating the true dynamics is an important research topic that has attracted considerable attentions in the past years; see [25, 54, 26, 8, 30] for studies on the approximation errors using conditional expectation [24]. Furthermore, efficient algorithms which allow for automatic identification of ”good” CVs are currently under rapid development; see for instance [57, 3, 51] for the recent progresses in finding CVs using machine learning and deep learning techniques.
@@
This paper is concerned with collective variables of discrete-in-time Markov processes $X_{n}$ in $\mathbb{R}^{d}$ in high dimension $d$, where $X_{n}$ may also be understood as resulting from a continuous-time Markov process by subsampling with constant lag-time. Collective variables are described by CV maps $\xi:\mathbb{R}^{d}\rightarrow\mathbb{R}^{k}$, where $1\leq k<d$, with an interest in $k\ll d$. We define the effective dynamics under a given CV map $\xi$ as the best Markovian representation of $X_{n}$ under $\xi$. The main goal is to analyze how good the effective dynamics is in approximating the original process $X_{n}$ on large timescales. For this purpose, we first lay out the theoretical basis by collecting results on how the largest (relaxation) timescales of $X_{n}$ are connected to the dominant eigenvalues of the transfer operator as well as how transition rates between metastable sets can be computed using the committor function via transition path theory (TPT) [46, 9, 10]. These well-known results are reformulated in a variational setting that is instrumental for the subsequent analysis.
@@
The novelty of the paper is that it gives strict criteria for selecting optimal collective variables via the properties of the effective dynamics. Concretely, we give three optimality statements: First, we show that the transition density of the effective dynamics of the optimal CV solves a relative entropy minimization problem in the sense that its Kullback-Leibler (KL) ”distance” to the transition density of $X_{n}$ is minimal, see Theorem 3.1. Next, we obtain quantitative error estimates for the effective dynamics in approximating the eigenvalues and transition rates of the original process. Based on these results, optimal CVs for minimizing these approximation errors are characterized, see Theorems 4.2 and 4.3.
@@
The remainder of this article is organized as follows. In Section 2, we study the transfer operator approach for analyzing metastable Markov processes. We present variational characterizations of the system’s main timescales via spectral analysis of the associated transfer operator, and of transition rates using TPT. In Section 3, we define the effective dynamics of discrete-in-time Markov processes in $\mathbb{R}^{d}$ for a given CV map and study its properties, in particular the relative entropy minimization property of the optimal CV. In Section 4, we analyze approximation errors of effective dynamics in terms of estimating eigenvalues and transition rates of the original process, and discuss algorithmic identification of optimal CVs.
@@ 2.1 Definitions and basic properties
Consider a stationary, discrete-in-time Markov process $X_{n}$ in $\mathbb{R}^{d}$ with a positive transition density function $p(x,y)$ for $x,y\in\mathbb{R}^{d}$, i.e. $\mathbb{P}(X_{n+1}\in D|X_{n}=x)=\int_{D}p(x,y)\,\mathrm{d}{y}$ for any measurable subset $D\subseteq\mathbb{R}^{d}$ and any integer $n\geq 0$. We assume that $X_{n}$ is ergodic with respect to a unique invariant distribution $\mu$, whose probability density $\pi$ is positive and satisfies $\mu(\mathrm{d}{x})=\pi(x)\mathrm{d}{x}$ and $\int_{\mathbb{R}^{d}}\pi(x)\,\mathrm{d}{x}=1$. In terms of density, the invariance of $\mu$ reads
$\int_{\mathbb{R}^{d}}p(x,y)\pi(x)\,\mathrm{d}{x}=\pi(y),\quad y\in\mathbb{R}^{d}\,.$ (1)
@@
where $\mathbb{E}(\cdot|X_{0}=x)$ denotes the expectation conditioned on the initial state $X_{0}=x$ (see Remark 2.1 below for discussions on different operators studied in the literature as well as their connections). It is natural to study $\mathcal{T}$ in the Hilbert space $\mathcal{H}=L^{2}_{\mu}(\mathbb{R}^{d})=\{f\,|\,f:\mathbb{R}^{d}\rightarrow\mathbb{R},\,\int_{\mathbb{R}^{d}}f(x)^{2}\pi(x)\,\mathrm{d}{x}<+\infty\}$ endowed with the weighted inner product (and the norm denoted by $\|\cdot\|_{\mu}$)
$\langle f,g\rangle_{\mu}=\int_{\mathbb{R}^{d}}f(x)g(x)\pi(x)\,\mathrm{d}{x}\,,\quad\forall f,g\in\mathcal{H}\,.$ (3)
With the above preparation we study the reversibility of $X_{n}$. Denote by $\mathcal{T}^{*}$ the adjoint of $\mathcal{T}$ in $\mathcal{H}$, such that $\langle\mathcal{T}f,g\rangle_{\mu}=\langle f,\mathcal{T}^{*}g\rangle_{\mu}$, for all $f,g\in\mathcal{H}$. It is straightforward to verify that
@@
Moreover, from (1) and (5) we can easily verify that $\int_{\mathbb{R}^{d}}p^{*}(x,y)\,\mathrm{d}{y}=1$ for all $x\in\mathbb{R}^{d}$ and $\int_{\mathbb{R}^{d}}p^{*}(x,y)\pi(x)\,\mathrm{d}{x}=\pi(y)$ for all $y\in\mathbb{R}^{d}$. Therefore, $p^{*}$ is a transition density function and it defines another Markov process (i.e. the adjoint of $X_{n}$) which has the same invariant distribution $\mu$.
@@
for test functions $f:\mathbb{R}^{d}\rightarrow\mathbb{R}$. From (4), (5), and the simple decomposition in (6)–(8), we obtain the following equivalent conditions which characterize the reversibility of the process $X_{n}$:
$\begin{split}&X_{n}~{}\mbox{satisfies detailed balance}\\ \Longleftrightarrow\quad&p^{*}\equiv p\\ \Longleftrightarrow\quad&\mathcal{T}=\mathcal{T}^{*}=\mathcal{T}^{rev}\\ \Longleftrightarrow\quad&\mathcal{T}^{non}=0\,.\end{split}$ (9)
Let us define the Dirichlet form [13]
$\mathcal{E}(f,g):=\frac{1}{2}\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d}}p(x,y)(f(y)-f(x))(g(y)-g(x))\pi(x)\,\mathrm{d}{x}\,\mathrm{d}{y}\,,$ (10)
for two test functions $f,g:\mathbb{R}^{d}\rightarrow\mathbb{R}$, and the Dirichlet energy
$\mathcal{E}(f):=\mathcal{E}(f,f)=\frac{1}{2}\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d}}p(x,y)(f(y)-f(x))^{2}\pi(x)\,\mathrm{d}{x}\,\mathrm{d}{y}\,.$ (11)
The following lemma summarizes the basic properties of these quantities.
@@
For any $f,g\in\mathcal{H}$, we have
$\mathcal{E}(f,g)=\langle(\mathcal{I}-\mathcal{T}^{rev})f,g\rangle_{\mu}\,,$ (12)
@@
$\mathcal{E}(f,g)=$ $\frac{1}{2}\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d}}p(x,y)\Big{(}f(y)g(y)+f(x)g(x)-f(y)g(x)-f(x)g(y)\Big{)}\pi(x)\,\mathrm{d}{x}\,\mathrm{d}{y}$
$=$ $\int_{\mathbb{R}^{d}}f(x)g(x)\pi(x)\,\mathrm{d}{x}-\frac{1}{2}\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d}}p(x,y)\Big{(}f(y)g(x)+f(x)g(y)\Big{)}\pi(x)\,\mathrm{d}{x}\,\mathrm{d}{y}$
$=$ $\int_{\mathbb{R}^{d}}f(x)g(x)\pi(x)\,\mathrm{d}{x}-\frac{1}{2}\Big{(}\langle\mathcal{T}f,g\rangle_{\mu}+\langle f,\mathcal{T}g\rangle_{\mu}\Big{)}$
$=$ $\langle(\mathcal{I}-\mathcal{T}^{rev})f,g\rangle_{\mu}$
where we used (1),(2) and (7) to obtain the second, the third and the last equality, respectively. The identity (13) follows directly from (11) and the ergodicity of the process. ∎
@@
We conclude this section with a discussion on the concrete examples that are mostly interesting to this work.
@@
$\mathrm{d}{x}(t)=b(x(t))\,\mathrm{d}{t}+\sigma(x(t))\,\mathrm{d}{w}(t)\,,\quad t\geq 0\,,$ (14)
with some drift and diffusion coefficients $b:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ and $\sigma:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d\times d}$, respectively, where $\tau>0$ is called the lag-time [39], $x(t)\in\mathbb{R}^{d}$ denotes the system’s state at time $t$, and $w(t)$ is a $d$-dimensional Brownian motion. In this case, the transition density $p(x,y)$ of $X_{n}$ is related to the transition density $p^{\mathrm{SDE}}(t,y\,|x)$ of $x(t)$ in (14) at time $t=\tau$, starting from $x\in\mathbb{R}^{d}$ at time $t=0$. The latter solves the Fokker-Planck equation [37, 41]
@@
for $t\in(0,+\infty)$ and $y\in\mathbb{R}^{d}$, where the operator $\mathcal{L}^{\top}$, given by $\mathcal{L}^{\top}f=-\mathrm{div}(bf)+\frac{1}{2}\sum_{i,j=1}^{d}\frac{\partial^{2}\big{(}(\sigma\sigma^{\top})_{ij}f\big{)}}{\partial x_{i}x_{j}}$ for a test function $f$, is the adjoint (with respect to the Lebesque measure) of the generator of (14), and $\delta(\cdot)$ denotes the Dirac delta function. In many cases, the solution to (15) is smooth for $t>0$ (see [37, Chapter 4] and [4] for a comprehensive study). In the following, we consider the transition density $p(x,y)$ in three concrete examples.
@@
    $\mathrm{d}{x}(t)=-\nabla V(x(t))\,\mathrm{d}{t}+\sqrt{2\beta^{-1}}\mathrm{d}{w}(t)\,,\quad t\geq 0$ (16)
    where $V:\mathbb{R}^{d}\rightarrow\mathbb{R}$ is a smooth potential function and $\beta>0$ is a constant related to the system’s temperature. The transition density of $X_{n}$ is then $p(x,y)=p^{\mathrm{BD}}(\tau,y\,|x)$, where $p^{\mathrm{BD}}(t,y\,|x)$ is the solution to (15) with the generator $\mathcal{L}f=-\nabla V\cdot\nabla f+\frac{1}{\beta}\Delta f$. Under certain conditions on $V$, $X_{n}$ is ergodic with the unique invariant distribution $\mu$ whose density is given by $\pi(x)=\frac{1}{Z}\mathrm{e}^{-\beta V(x)}$, where $Z=\int_{\mathbb{R}^{d}}\mathrm{e}^{-\beta V(x)}\,\mathrm{d}{x}$ is a normalizing constant. In particular, $X_{n}$ is reversible and the corresponding transfer operator defined in (2) is given by $\mathcal{T}f=\mathrm{e}^{\tau\mathcal{L}}f$ (the solution to the backward Kolmogorov equation associated to $\mathcal{L}$ [36, Chapter 8]).
@@
    Langevin dynamics. Assume that the evolution of $x(s)$ is governed by the (underdamped) Langevin dynamics [28, Section 2.2.3]
    $\begin{split}\mathrm{d}{x}(t)=&v(t)\mathrm{d}{t}\\ \mathrm{d}{v}(t)=&-\nabla V(x(t))\mathrm{d}{t}-\gamma\,v(t)\mathrm{d}{t}+\sqrt{2\gamma\beta^{-1}}\mathrm{d}{w}(t)\end{split}$ (17)
    where $v(t)\in\mathbb{R}^{d}$ denotes the system’s velocity and $\gamma>0$ is a friction constant. In this case, if we make the assumption that the (position) data $X_{0},X_{1},X_{2},\dots$ comes from a Markovian process (which is not true) and estimate its transition density using a density estimation method, then the ideal transition density (obtained as data size goes to infinite) is given by
    $p(x,y)=(\frac{\beta}{2\pi})^{\frac{d}{2}}\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d}}p^{\mathrm{Lan}}(\tau,y,v^{\prime}|x,v)\,\mathrm{e}^{-\frac{\beta|v|^{2}}{2}}\,\mathrm{d}{v}\,\mathrm{d}{v^{\prime}},$
    where $p^{\mathrm{Lan}}(\tau,y,v^{\prime}|x,v)$ denotes the transition density of (17). We refer to Appendix A for detailed derivations.
@@
    where $(W_{n})_{n\geq 0}$ are independent standard Gaussian variables in $\mathbb{R}^{d}$ and $\Delta t$ is the step-size (here, for simplicity we assume Euler-Maruyama scheme [19, Chapter 9] and $\tau=\Delta t$). In this case, the transition density of $X_{n}$ is given by $p(x,y)=\big{(}\frac{\beta}{4\pi\Delta t}\big{)}^{\frac{d}{2}}\exp(-\frac{\beta|y-x+\nabla V(x)\Delta t|^{2}}{4\Delta t})$, for $x,y\in\mathbb{R}^{d}$. The process $X_{n}$ is in general not reversible, but in practice this irreversibility due to numerical discretization is often ignored whenever $\Delta t$ is small. Alternatively, reversibility can be retained by adding a Matropolis-Hasting acceptance/rejection step [28, Chapter 2].

### v2

@@ 1 Introduction
A commonly adopted strategy is to utilize the fact that the dynamics of high-dimensional systems, e.g. molecular systems in MD and materials science, can often be well characterized using only a few observables, often called collective variables (CVs) or reaction coordinates, of the systems. In particular, many enhanced sampling methods (see [22] for a review) utilize the knowledge about the system’s CVs to perform accelerated sampling and efficient free energy calculations [37], and various methods for model reduction rely on this knowledge to construct simpler (lower-dimensional) surrogate models [44, 33, 3]. While these methods have proven to be greatly helpful in understanding dynamical behaviors of complex systems, the accuracy of estimated quantities and surrogate models offered by these methods crucially depend on the choices of CVs. Given a set of CVs, analyzing the accuracy of surrogate models in approximating the true dynamics is an important research topic that has attracted considerable attentions in the past years; see [33, 34, 66, 35, 14, 39] for studies on effective dynamics constructed using conditional expectation. These work extend the classical averaging technique for multiscale processes to the setting where the slow variable is nonlinear (see [47, Chapter 10] for averaging techniques and [39, 35] for discussions on the connection between averaging and effective dynamics by conditional expectation). Furthermore, efficient algorithms which allow for automatic identification of “good” CVs are currently under rapid development; see for instance [69, 7, 62] for recent progresses in finding CVs using machine learning and deep learning techniques.
@@
This paper is concerned with collective variables of discrete-in-time Markov processes $X_{n}$ in $\mathbb{R}^{d}$ in high dimension $d$, where $X_{n}$ may also be understood as resulting from a continuous-time Markov process by subsampling with constant lag-time. Collective variables are described by CV maps $\xi:\mathbb{R}^{d}\rightarrow\mathbb{R}^{k}$, where $1\leq k<d$, with an interest in $k\ll d$. Following the idea in [33], we define the effective dynamics of $X_{n}$ under a given CV map $\xi$ using conditional expectations. The main goal is to analyze how good the effective dynamics is in approximating the original process $X_{n}$ on large timescales, and also use this analysis to propose numerical approaches for optimizing CV maps. For this purpose, we first lay out the theoretical basis (in Section 2) by collecting results on how the largest (relaxation) timescales of $X_{n}$ are connected to the dominant eigenvalues of the transfer operator, as well as how transition rates between metastable sets can be computed using the committor function via transition path theory (TPT) [57, 15, 16]. These well-known results are reformulated in a variational setting that is instrumental for the subsequent analysis.
@@
The novelty of the paper is that it gives strict criteria for selecting optimal collective variables via the properties of the effective dynamics. Concretely, we give three optimality statements: First, we show that the effective dynamics introduced in this paper (Definition 43) is the optimal surrogate model in the sense that the Kullback-Leibler (KL) “distance” to the transition density of $X_{n}$ is minimized (Proposition 6). This optimality allows us to define the optimal CV map as the solution to a relative entropy minimization problem (Theorem 3.1). Next, we obtain quantitative error estimates for the effective dynamics in approximating the eigenvalues and transition rates of the original process. Based on these results, optimal CVs for minimizing these approximation errors are characterized, see Theorems 4.2 and 4.3.
@@
The remainder of this article is organized as follows. In Section 2, we study the transfer operator approach for analyzing metastable Markov processes. We present variational characterizations of the system’s main timescales via spectral analysis of the associated transfer operator, and of transition rates using TPT. In Section 3, we define the effective dynamics of discrete-in-time Markov processes for a given CV map and study its properties, in particular the relative entropy minimization property of the optimal CV. In Section 4, we analyze approximation errors of effective dynamics in terms of estimating eigenvalues and transition rates of the original process, and discuss algorithmic identification of optimal CVs. In Appendix A, we provide details for the concrete case where the transfer operator models the evolution of position variables of Langevin dynamics. In Appendix B, we present the proofs of some results in Section 2.
@@ 2.1 Definitions and basic properties
Consider a stationary, discrete-in-time Markov process $X_{n}$ in $\mathbb{R}^{d}$ with a positive transition density function $p(x,y)$ for $x,y\in\mathbb{R}^{d}$, i.e. $\mathbb{P}(X_{n+1}\in D|X_{n}=x)=\int_{D}p(x,y)\,\mathrm{d}{y}$ for any measurable subset $D\subseteq\mathbb{R}^{d}$ and any integer $n\geq 0$. We assume that $X_{n}$ is ergodic with respect to a unique invariant distribution $\mu(\mathrm{d}{x})=\pi(x)\mathrm{d}{x}$, where the probability density $\pi$ is positive (see [31, Chapter 1] and [2, Section 2] for ergodicity of Markov processes). The invariance of $\mu$ means
$\left(\int_{\mathbb{R}^{d}}p(x,y)\mu(\mathrm{d}{x})\right)\mathrm{d}{y}=\mu(\mathrm{d}{y}),~{}\mbox{or equivalently,}~{}\int_{\mathbb{R}^{d}}p(x,y)\mu(\mathrm{d}{x})=\pi(y),\quad y\in\mathbb{R}^{d}\,.$ (1)
@@
where $\mathbb{E}(\cdot|X_{0}=x)$ denotes the expectation conditioned on the initial state $X_{0}=x$ (see Remark 2.1 below for discussions on different operators studied in the literature as well as their connections). It is natural to study $\mathcal{T}$ in the Hilbert space
$\mathcal{H}=L^{2}_{\mu}(\mathbb{R}^{d})=\left\{f\,\middle|\,f:\mathbb{R}^{d}\rightarrow\mathbb{R},\,\int_{\mathbb{R}^{d}}f^{2}(x)\mu(\mathrm{d}{x})<+\infty\right\}$
endowed with the weighted inner product (and the norm denoted by $\|\cdot\|_{\mu}$)
$\langle f,h\rangle_{\mu}=\int_{\mathbb{R}^{d}}f(x)h(x)\mu(\mathrm{d}{x})\,,\quad\forall f,h\in\mathcal{H}\,.$ (3)
With the above preparation we study the reversibility of $X_{n}$. Denote by $\mathcal{T}^{*}$ the adjoint of $\mathcal{T}$ in $\mathcal{H}$, such that $\langle\mathcal{T}f,h\rangle_{\mu}=\langle f,\mathcal{T}^{*}h\rangle_{\mu}$, for all $f,h\in\mathcal{H}$. It is straightforward to verify that
@@
Moreover, from (1) and (5) we can easily verify that $\int_{\mathbb{R}^{d}}p^{*}(x,y)\,\mathrm{d}{y}=1$ for all $x\in\mathbb{R}^{d}$ and $\int_{\mathbb{R}^{d}}p^{*}(x,y)\,\mu(\mathrm{d}{x})=\pi(y)$ for all $y\in\mathbb{R}^{d}$. Therefore, $p^{*}$ is a transition density function and it defines another Markov process (i.e. the adjoint of $X_{n}$) which has the same invariant distribution $\mu$.
@@
for test functions $f:\mathbb{R}^{d}\rightarrow\mathbb{R}$. From (4), (5), and the simple decomposition in (6)–(8), we obtain the following equivalent conditions.
$\begin{split}&~{}p^{*}\equiv p\hskip 56.9055pt\mbox{(detailed balance)}\\ \Longleftrightarrow&~{}\mathcal{T}=\mathcal{T}^{*}=\mathcal{T}^{rev}\hskip 17.07182pt\mbox{(self-adjointness)}\\ \Longleftrightarrow&~{}\mathcal{T}^{non}=0\,.\end{split}$ (9)
We will refer to (9) as reversibility conditions and say that $X_{n}$ is reversible if any of the equivalent conditions in (9) is satisfied.
Let us define the Dirichlet form [19, Chapter 1]
$\mathcal{E}(f,h):=\frac{1}{2}\int_{\mathbb{R}^{d}}\left(\int_{\mathbb{R}^{d}}(f(y)-f(x))(h(y)-h(x))p(x,y)\mathrm{d}{y}\,\right)\mu(\mathrm{d}{x})\,,$ (10)
for two test functions $f,h:\mathbb{R}^{d}\rightarrow\mathbb{R}$, and the Dirichlet energy
$\mathcal{E}(f):=\mathcal{E}(f,f)=\frac{1}{2}\int_{\mathbb{R}^{d}}\left(\int_{\mathbb{R}^{d}}(f(y)-f(x))^{2}p(x,y)\mathrm{d}{y}\,\right)\mu(\mathrm{d}{x})\,.$ (11)
The following lemma summarizes basic properties of these quantities.
@@
For any $f,h\in\mathcal{H}$, we have
$\mathcal{E}(f,h)=\langle(\mathcal{I}-\mathcal{T}^{rev})f,h\rangle_{\mu}\,,$ (12)
@@
$\mathcal{E}(f,h)=$ $\frac{1}{2}\int_{\mathbb{R}^{d}}\left(\int_{\mathbb{R}^{d}}\Big{(}f(y)h(y)+f(x)h(x)-f(y)h(x)-f(x)h(y)\Big{)}\,p(x,y)\mathrm{d}{y}\right)\mu(\mathrm{d}{x})$
$=$ $\int_{\mathbb{R}^{d}}f(x)h(x)\,\mu(\mathrm{d}{x})-\frac{1}{2}\int_{\mathbb{R}^{d}}\left(\int_{\mathbb{R}^{d}}\Big{(}f(y)h(x)+f(x)h(y)\Big{)}p(x,y)\mathrm{d}{y}\right)\mu(\mathrm{d}{x})$
$=$ $\int_{\mathbb{R}^{d}}f(x)h(x)\,\mu(\mathrm{d}{x})-\frac{1}{2}\Big{(}\langle\mathcal{T}f,h\rangle_{\mu}+\langle f,\mathcal{T}h\rangle_{\mu}\Big{)}$
$=$ $\langle(\mathcal{I}-\mathcal{T}^{rev})f,h\rangle_{\mu}$
where we used (1),(2) and (7) to obtain the second, the third and the last equality, respectively. Since we assume that the process is ergodic, the identity (13) follows directly from (11) and Birkhoff’s ergodic theorem (see [31, Theorem 4.4 in Chapter 1]). ∎
@@
We conclude this section with a discussion on the concrete examples that are most interesting to this work.
@@
$\mathrm{d}{x}(t)=b(x(t))\,\mathrm{d}{t}+\sigma(x(t))\,\mathrm{d}{\eta}(t)\,,\quad t\geq 0\,,$ (14)
with some drift and diffusion coefficients $b:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ and $\sigma:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d\times d}$, respectively, where $\tau>0$ is called the lag-time [50], $x(t)\in\mathbb{R}^{d}$ denotes the system’s state at time $t$, and $\eta(t)$ is a $d$-dimensional Brownian motion. In this case, the transition density $p(x,y)$ of $X_{n}$ is related to the transition density $p^{\mathrm{SDE}}(t,y\,|x)$ of $x(t)$ in (14) at time $t=\tau$, starting from $x\in\mathbb{R}^{d}$ at time $t=0$. The latter solves the Fokker-Planck equation [48, 52]
@@
for $t\in(0,+\infty)$ and $y\in\mathbb{R}^{d}$, where the operator $\mathcal{L}^{\top}$, given by $\mathcal{L}^{\top}f=-\mathrm{div}(bf)+\frac{1}{2}\sum_{i,j=1}^{d}\frac{\partial^{2}\big{(}(\sigma\sigma^{\top})_{ij}f\big{)}}{\partial x_{i}x_{j}}$ for a test function $f$, is the adjoint (with respect to the Lebesgue measure) of the generator of (14), and $\delta(\cdot)$ denotes the Dirac delta function. In many cases, the solution to (15) is smooth for $t>0$ (see [48, Chapter 4] and [8] for a comprehensive study). In the following, we consider the transition density $p(x,y)$ in three concrete examples.
@@
    $\mathrm{d}{x}(t)=-\nabla V(x(t))\,\mathrm{d}{t}+\sqrt{2\beta^{-1}}\mathrm{d}{\eta}(t)\,,\quad t\geq 0$ (16)
    where $V:\mathbb{R}^{d}\rightarrow\mathbb{R}$ is a smooth potential function and $\beta>0$ is a constant related to the system’s temperature. The transition density of $X_{n}$ is then $p(x,y)=p^{\mathrm{BD}}(\tau,y\,|x)$, where $p^{\mathrm{BD}}(t,y\,|x)$ is the solution to (15) with the generator $\mathcal{L}f=-\nabla V\cdot\nabla f+\frac{1}{\beta}\Delta f$. Under certain conditions on $V$, $X_{n}$ is ergodic with the unique invariant distribution $\mu$ whose density is given by $\pi(x)=\frac{1}{Z}\mathrm{e}^{-\beta V(x)}$, where $Z=\int_{\mathbb{R}^{d}}\mathrm{e}^{-\beta V(x)}\,\mathrm{d}{x}$ is a normalizing constant. In particular, $X_{n}$ is reversible and the corresponding transfer operator defined in (2) is given by $\mathcal{T}f=\mathrm{e}^{\tau\mathcal{L}}f$ (the solution to the backward Kolmogorov equation associated to $\mathcal{L}$ [46, Chapter 8]).
@@
    Langevin dynamics. Assume that the evolution of $x(s)$ is governed by the (underdamped) Langevin dynamics [37, Section 2.2.3]
    $\begin{split}\mathrm{d}{x}(t)=&v(t)\mathrm{d}{t}\\ \mathrm{d}{v}(t)=&-\nabla V(x(t))\mathrm{d}{t}-\gamma\,v(t)\mathrm{d}{t}+\sqrt{2\gamma\beta^{-1}}\mathrm{d}{\eta}(t)\end{split}$ (17)
    where $v(t)\in\mathbb{R}^{d}$ denotes the system’s velocity and $\gamma>0$ is a friction constant. In this case, if we make the assumption that the (position) data $X_{0},X_{1},X_{2},\dots$ comes from a Markov process and estimate its transition density using a density estimation method, then the ideal transition density (obtained as data size goes to infinite) is given by
    $p(x,y)=(\frac{\beta}{2\pi})^{\frac{d}{2}}\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d}}p^{\mathrm{Lan}}(\tau,y,v^{\prime}|x,v)\,\mathrm{e}^{-\frac{\beta|v|^{2}}{2}}\,\mathrm{d}{v}\,\mathrm{d}{v^{\prime}}\,,$ (18)
    where $p^{\mathrm{Lan}}(\tau,y,v^{\prime}|x,v)$ denotes the transition density of (17). We refer to Appendix A for detailed derivations. In particular, we show that the Markov process described by $p(x,y)$ in (18) is reversible, although the original Langevin dynamics is non-reversible.
@@
    where $(W_{n})_{n\geq 0}$ are independent standard Gaussian variables in $\mathbb{R}^{d}$ and $\Delta t$ is the step-size (here, for simplicity we assume Euler-Maruyama scheme [26, Chapter 9] and $\tau=\Delta t$). In this case, the transition density of $X_{n}$ is given by $p(x,y)=\big{(}\frac{\beta}{4\pi\Delta t}\big{)}^{\frac{d}{2}}\exp(-\frac{\beta|y-x+\nabla V(x)\Delta t|^{2}}{4\Delta t})$, for $x,y\in\mathbb{R}^{d}$. The process $X_{n}$ is in general not reversible, but in practice this irreversibility due to numerical discretization is often ignored whenever $\Delta t$ is small. Alternatively, reversibility can be retained by adding a Matropolis-Hasting acceptance/rejection step [37, Chapter 2].